OpenStudy (anonymous):

find critical numbers for x^3-6x^2+15

OpenStudy (johnweldon1993):

Critical Numbers are when the derivative = 0 right? so what is the derivative of this function?

OpenStudy (anonymous):

yes. so i got 3x^2-12x=0

OpenStudy (johnweldon1993):

Alright good...so we have 3x² - 12x = 0 we can factor out a 3 of this equation here... 3(x² - 4x) can we factor x² - 4x ?

OpenStudy (anonymous):

i had 3(x^2-4x) so to factor x^2-4x would it be x(x-4)?

OpenStudy (johnweldon1993):

That would be correct...so what values for 'x' would make 3x(x-4) = 0 ? well...either 3x = 0 or x - 4 = 0 so what 2 values do we have?

OpenStudy (anonymous):

would 4 be one ?

OpenStudy (johnweldon1993):

yes 4 would be one value... what about the 3x = 0 ? what 'x' would make this = 0?

OpenStudy (anonymous):

just 0 ?

OpenStudy (johnweldon1993):

mmhmm :) and those are the only 2 critical points of this graph!

OpenStudy (anonymous):

ahh thank you! now when it says find the open intervals this is increasing or decreasing would it be negative infinity to 0 and 5to infinity ?

OpenStudy (johnweldon1993):

Well using correct notation...I believe it should be [-infinity , 0) && (4 , infinity] [ and ] means the number is included ( and ) means the number is NOT included

OpenStudy (anonymous):

yes im sorry, i meant 4 not 5. and then the relative max would be 4? and the relative min would be 0?

OpenStudy (johnweldon1993):

Not quite......Judging from the equation your graph will most likely look something like...  bad drawing...but regardless...relative max would be? and relative min would be?